Answer=25%
14 red marbles
12 yellow marbles
26 total marbles
To solve this, we need to first find out the probability of the individual events...
The odds of selecting a red marble is 14 out of 26 total marbles: 14/26
The odds of selecting a yellow marble after the first marble was NOT replaced is 12 out of 25. (25 since the first marble wasn't replaced: 26-1=25) 12/25
Now lets multiple both probabilities together to solve for the chance of both happening consecutively....
14/26*12/25= 168/650=84/325
84/325=0.25846153846= 25%